divide the product of 1/2 and 2/11 by 3 1/2
Found 2 solutions by edjones, cel072880:
Answer by edjones(7311)

1/2 * 2/11 = 2/22 = 1/11
1/11 / 7/2
=1/11 * 2/7
=2/77
.
Ed

Answer by cel072880(4)

First, we will multiply 1/2 and 2/11
1/2 * 2/11= 2/22 or 1/11 in lowest term
Then we will convert 3 1/2 to improper fraction
3 1/2 = 7/2
Then, we will get the quotient of 1/11 and 7/2
1/11 / 7/2
= 1/11 * by the reciprocal of 7/2
= 1/11 * 2/7 = 2/77
Therefore, the answer is 2/77